Assume that the causes of a disease are 29% of the time associated with diet and 71% of the time associated with genetics, and that the causes of the disease are independent from each other. What is the expected number of cases of the disease due to genetics before the first case appears due to diet?
Answer #1

from geometric distribution

expected number of cases of the disease due to genetics before the first case appears due to diet =1/p-1=1/0.29-1=2.4483